OpenStudy (anonymous):

Lane has a gift card worth $300 for a local entertainment store. Movies cost $20 each and newly released video games cost $50 each. In order to use the gift card, Lane must purchase at least seven items. Which is a combination of movies and newly released video games that Lane could purchase using his gift card?

OpenStudy (anonymous):

my answer choices are 1 movie and 6 video games 2 movies and 4 video games 6 movies and 5 video games 8 movies and 1 video game

OpenStudy (anonymous):

let there be x movies and y video games. so 20x+50y=300 2x+5y=30 when you plot the graph for this eq. it is a line passing through (15,0) and (0,6)... x should be divisible by 5 and y should be divisible by 2. so facing all these constraints the possible solutions are x=0 y=6 x=5 y=4 x=10 y=2 x=15 y=0 but x+y >7 so x=0 y=6 is not valid

OpenStudy (anonymous):

so are you saying the amount of video games and movies is less than 7?

OpenStudy (anonymous):

no... i mean either x=5 y=4 or x=10 y=2 or x=15 y=0

OpenStudy (anonymous):

that doesnt match my answer choices

OpenStudy (anonymous):

yes... these do not match the options you gave :(